An engine cylinder has a cross-sectional area of 0.010 m. How much work can be done by a gas
in the cylinder if the gas exerts a constant pressure of 7.5 × 105 Pa on the piston and moves the
piston a distance of 0.040 m?
1
Expert's answer
2011-08-25T13:29:06-0400
Let S be the area of a cross-section of the cylinder, p - pressure of the gas, d = 0.04 m.
The total work can be calculated as:
A = F*d.

The force F can be found from the equation:

p = F/S, => F=p*S.

So:

A = p*S*d.

A = 7.5*105*0.01*0.04 = 300 J

So, the gas can do 300 J of work.